OpenStudy (anonymous):

integral of (2x+1)(x^2+x)^9 dx, using u=x^2+x

OpenStudy (anonymous):

If u = x^2 + x, then du = (2x+1) dx So the integral is of the form, u^9 du and the integral of u^9 du is just u^10/10 So substitute u = 2x + 1 for u, so your final answer is (2x+1)610/10 + c Dont forget the + C at the end.

OpenStudy (anonymous):

is that a 610 or ^10?

OpenStudy (anonymous):

ir recon it is, thanks mate i got it!!!

OpenStudy (anonymous):

Sorry, thats(2x+1)^10

OpenStudy (anonymous):

typo error

OpenStudy (anonymous):

awesome, do you mind if i ask another one that has got sim and cos in ti???

OpenStudy (anonymous):

***sin and cos in it???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

integral of (sinx)/(cos^4x)dx, using u = cos x

OpenStudy (anonymous):

let u= cos (x); so du = - sin x dx So what form does this integral take? Thats the crucial question, the form that the integral takes. Well......

OpenStudy (anonymous):

it just tells me to solve the integral but i have to use that specific substitution to do it

OpenStudy (anonymous):

We pull the minus sign in front of the integral, and it takes the form of u^4 du Agree?????????

OpenStudy (anonymous):

Let me ne very specific...the integral is (sin x) (cos^4x) ds....so the sin x dx...thats just -du (because du = - sin x)

OpenStudy (anonymous):

So the inegrand (whats inside the integral sin) is of the form u^4 du

OpenStudy (anonymous):

agree?

OpenStudy (anonymous):

i agree that its -du

OpenStudy (anonymous):

If u = cos x and du = - sin x dx By substitution, isnt (sin x)(cos^4x) dx = - u^5 du??????

OpenStudy (anonymous):

sorry, I meant u^4 du

OpenStudy (anonymous):

and just be clear only the 4 is ^ in the cos part not the x does that make sense to you

OpenStudy (anonymous):

u = cos x , so cos^4x is just u^4

OpenStudy (anonymous):

correct..only the 4 is in the exponent

OpenStudy (anonymous):

cool thats the part that was messing with my head lol

OpenStudy (anonymous):

so the integral is of the form of u^4 du with a minus sign on the outside of the integral

OpenStudy (anonymous):

The integral of u^4 du is just u^5/5 + C...agree???????

OpenStudy (anonymous):

yep that sounds right and then all i have to do is substitute cos x back in right?

OpenStudy (anonymous):

exactly...final answer is - (cos x)^5/5 + C Dont forget that minus sign in the front as we pulled out the minus sign

OpenStudy (anonymous):

ohh right. thanks mate your a real life legend. it was stumping me for ages

OpenStudy (anonymous):

Review this problem..I dont want it to sound right..I want to make sure that you are crystal clear about this problem!

OpenStudy (anonymous):

alright hang on an ill read through it. i sec

OpenStudy (anonymous):

It's not a novel for reading..It must be gone through very carefully with a solid understanding!

OpenStudy (anonymous):

yep, i get it thats heaps great. the 4 in between the cos and x was getting to me but i understand now, thanks heaps